# These $\altaffiliationThese authors contributed equally.$

### v1

@@ Abstract
Analyzing nonadiabatic molecular dynamics trajectories traditionally heavily relies on expert intuition and visual pattern recognition, a process that is difficult to formalize. We present VisU, a vision-driven framework that leverages the complementary strengths of two state-of-the-art large language models to establish a “virtual research collective.” This collective operates through a “Mentor-Engineer-Student” paradigm that mimics the collaborative intelligence of a professional chemistry laboratory. Within this ecosystem, the Mentor provides physical intuition through visual reasoning, while the Engineer adaptively constructs analysis scripts, and the Student executes the pipeline and managing the data and results. VisU autonomously orchestrates a four-stage workflow comprising Preprocessing, Recursive Channel Discovery, Important-Motion Identification, and Validation/Summary. This systematic approach identifies reaction channels and key nuclear motions while generating professional academic reports. By bridging visual insight with chemical expertise, VisU establishes a new paradigm for human-AI collaboration in the analysis of excited-state dynamics simulation results, significantly reducing dependence on manual interpretation and enabling more intuitive, scalable mechanistic discovery.
@@ 1 Introduction
Extracting chemically meaningful insights from the output data obtained from the trajectory-based NAMD simulations requires the integration of visual reasoning with domain-specific knowledge. Researchers routinely inspect trajectory evolutions, critical structures and dynamical animations to correlate molecular spatial patterns with photochemical concepts such as conical intersections and reaction coordinates. This tightly-coupled visual-chemical reasoning loop is essential for distinguishing nonadiabatic channels and identifying governing nuclear motions.
Recent advances in unsupervised machine learning (ML) provided valuable tools for dimensionality reduction and clustering to handle high-dimensional trajectory data. zhuUnsupervisedMachineLearning2024, zhu2022analysis, glielmo2021unsupervised, ceriottiUnsupervisedMachineLearning2019, virshup2012nonlinear Nevertheless, these methods operate in a perceptual vacuum. They lack the ability to perceive data through the lens of a chemist and cannot integrate visual patterns with domain knowledge to evaluate the chemical significance of identified clusters. Consequently, advanced automated pipelines zhuPrincipalComponentAnalysis2022a, zhuUnsupervisedMachineLearning2024, pinheiroULaMDynEnhancingExcitedstate2025, liu2025automatedframeworkanalyzingstructural still require continuous human intervention for visual validation and parameter tuning. zhuUnsupervisedMachineLearning2024 In one word, the human experts still play essential roles in such analysis protocols, due to their visiual preception and professional knowledge.
The rise of large language models (LLMs) offers new opportunities for orchestrating scientific workflows acrossing various scientific research fields. ai4science2023impactlargelanguagemodels, jimenez2024swebench, laurent2024labbenchmeasuringcapabilitieslanguage, miret2025enabling, white2023future, jablonka202314, ramos2025review, chenSciToolAgentKnowledgeGraphDriven2025 However, conventional LLM-based approaches remain poorly aligned with the demands of NAMD analysis. li2025chemvlmexploringpowermultimodal, chenLargescaleChemicalReaction2025 Most existing working paradigms apply LLMs as either autonomous planners or passive executors. Neither role is suitable for trajectory interpretation, which requires simultaneous chemical reasoning and visual comprehension of complex spatial patterns. Critically, text-only LLMs lack access to the visual cues essential for understanding molecular geometries and reaction pathways, leaving this domain largely unexplored.
@@
By orchestrating vision-capable LLMs within this structured framework, VisU operationalizes a perception-to-insight pipeline that bridges the gap between raw numerical data and interpretable chemical knowledge. In this framework, visual reasoning is not merely a supplementary tool but the primary driver of the iterative process, mimicking how a scientist’s eye discerns complex patterns to refine hypotheses and validate mechanisms. This work demonstrates that embedding autonomous visual intelligence into the NAMD analysis workflow redefines how we process excited-state dynamics. At the same time, this novel research paradigm offers a scalable template for human-AI collaboration across theoretical chemistry, which enables researchers to uncover molecular mechanisms from massive datasets with unprecedented clarity and efficiency.
@@ 2.1 A Brief Description of VisU
In this section, we provide an overview of VisU, an autonomous vision-AI-powered unsupervised analysis framework. VisU framework adopts a “Mentor-Engineer-Student” architecture, faithfully simulating the division of labor in a real nonadiabatic dynamics research group. In a traditional theoretical, three roles perform distinct but complementary functions in a joint project: an experienced mentor supervises the overall analysis and provides expert chemical guidance, a skilled programming-engineer handles computational tool construction and processing script orchestration, and a diligent student executes computational tasks while generating intermediate results. Under this human-cooperative paradigm, a comprehensive NAMD analysis typically proceeds through four successive stages: data preprocessing, unsupervised channel discovery, identification of dominant molecular motions, as well as validation and summarization.
VisU operationalizes this collaborative three-role, four-step paradigm into a fully automated end-to-end protocol:
@@
To demonstrate its potential, the current implementation of VisU builds upon the validated framework of previous work. zhuPrincipalComponentAnalysis2022a That work established a hierarchical analysis protocol for characterizing ring deformations in atomic systems via a combination of principal component analysis (PCA) wold1987principal, abdi2010principal and various clustering algorithms. Future iterations will expand the framework’s applicability to a broader spectrum of molecular architectures and complex photophysical processes.
@@
[H] [Uncaptioned image] Workflow of VisU, an autonomous vision-AI framework for hierarchical analysis of nonadiabatic dynamics. The system consists of four modules: (a) Preprocessing module autonomously parses raw NAMD outputs and generates standardized descriptor sets; (b) Recursive Channel Discovery module iteratively applies PCA-guided clustering to uncover chemically distinct reaction channels; (c) Important-Motion Identification module analyzes PCA projections and visualizes dominant nuclear displacements for each channel; (d) Validation and Mechanistic Summary module selects representative structures, performs final quality checks, and generates concise mechanistic interpretations.
@@ 2.1.1 Data Preprocessing
VisU accepts two input formats: (i) three mandatory XYZ files, that is, ref.xyz (reference geometry), hopping.xyz (all hopping geometries from NAMD trajectories), and init.xyz (corresponding initial geometries), or (ii) a single ZIP archive containing raw simulation outputs from any NAMD package. Although both approaches are allowed, the direct submission of the three XYZ files is highly recommended for routine analyses, as it saves tokens and eliminates any remaining risk of extraction errors.
@@
In general, different NAMD packages produce distinct output formats and directory structures. Conventional LLM-based workflow lack direct access to file contents and system details, preventing them from reliably interpreting such heterogeneous file systems during preprocessing. Without knowledge of the file contents or structures, these pipelines typically rely on multi-step trial-and-error workflows, iteratively attempting to interpret outputs, generate scripts, execute them, and refine the logic based on error messages. This process not only incurs high computational costs and carries a substantial risk of non-convergence but also requires extensive elaborate instruction frameworks to handle diverse scenarios.
@@
To ensure broad compatibility across the heterogeneous and often inconsistent output file formats generated by different NAMD simulation packages, we deliberately avoids allowing LLMs to generate a complete preprocessing script from scratch. Instead, the framework provides a curated library of expert-validated Python templates for structure extraction, format sanitization, and data integrity checks. With the support of these templates, the Engineer constructs the full preprocessing script by assembling and adapting predefined components rather than writing arbitrary code. This design substantially improves both flexibility and reliability in data preprocessing, surpassing conventional LangChain-style agents that are often constrained by static tool graphs or rigid, hard-coded workflows. Following script generation, the Student executes the preprocessing codes, generates all required files, and returns execution feedback. If execution fails, the complete traceback information is automatically fed back to the LLM, enabling iterative self-repair and refinement of the preprocessing operations until valid output files are produced. Overall, this template-grounded and self-repairing strategy establishes a more effective paradigm for automated scientific code generation and robust preprocessing of heterogeneous simulation outputs. Crucially, the tools and templates developed for the Engineer are designed to be general-purpose and system-agnostic, rather than being tailored to specific case studies. More details are provided in the SI.
@@
Following preprocessing, VisU proceeds to systematic structure analysis and descriptor generation through the close cooperation of the Mentor and Student. The core of this stage is the introduction of a dedicated vision-language chemistry Mentor, a large multimodal model extensively prompted with expert-level chemical knowledge. The workflow starts with the Student generating a high-quality 2D molecular diagram from the optimized reference geometry in ref.xyz.
@@
As the currently supported analysis workflow in VisU zhuPrincipalComponentAnalysis2022a primarily targets the geometric evolution of atomic ring systems in NAMD simulations, the following discussion focuses on such cases. This scope is sufficient to illustrate the effectiveness of the vision-powered “Mentor-Engineer-Student” cooperative strategy for NAMD analysis, while extensions to more general molecular topologies are left for future work.
Previous works zhuPrincipalComponentAnalysis2022a, liu2025automatedframeworkanalyzingstructural demonstrated that redundant internal coordinates (RICs) provide an appropriate molecular descriptor for NAMD analysis and are therefore adopted here. RICs consist of bond lengths (R), bond angles (A), and dihedral angles (D), which span different numerical ranges and encode distinct physical aspects of molecular structure. For systems containing atomic rings, RICs associated with the ring framework and those involving terminal groups play fundamentally different roles during nonadiabatic dynamics. Accordingly, within the workflow, the Mentor instructs the Student to classify atoms into ring (ring) and end-group (eg) subsets. The Student then constructs six descriptor groups, $R_{\mathrm{ring}}$, $A_{\mathrm{ring}}$, $D_{\mathrm{ring}}$, $R_{\mathrm{eg}}$, $A_{\mathrm{eg}}$, and $D_{\mathrm{eg}}$, which are subsequently used for unsupervised analysis. To avoid overemphasizing large-amplitude hydrogen motions, hydrogen atoms are excluded, and the analysis focuses exclusively on the molecular scaffold.
@@ 2.1.2 Recursive Channel Discovery
The selection of an appropriate descriptor subset follows a standardized visual screening procedure. The Student generates six low-dimensional projections corresponding to all descriptor groups and submits them to the Mentor for inspection. Based on the separability, compactness, and overall structure of the data distributions, the Mentor identifies the reduced spaces that exhibit the most distinguishable channel patterns and selects the corresponding reduced space for subsequent clustering analyses.
@@ 2.1.3 Important-Motion Identification
The identification of these critical motions is performed through a tight coordination between the Mentor and the Student. For each identified channel, the Student assembles the initial and hopping geometries, executes PCA, and generates a comprehensive visual package. This package includes 2D projection plots, heatmaps of the principal components, and three-view (front, side, and top) renderings of both the reference and averaged hopping structures. Under the guidance of its vision-reasoning capabilities, the Mentor evaluates these multi-faceted figures to pinpoint the active coordinates and leading molecular motions characteristic of the channel. This Mentor-Student joint procedure effectively bridges the gap between unsupervised data clustering and human-interpretable chemical mechanisms, establishing a fully automated NAMD analysis protocol that eliminates the requirement for manual validation.
@@ 2.1.4 Validation and Mechanistic Summary
For each reaction channel, the Student randomly samples a representative set of hopping geometries (typically ten) and prepares their three-view renderings for the Mentor. Upon visual evaluation, the Mentor identifies the most characteristic geometry based on previously established analysis results. Simultaneously, the Mentor inspects representative geometries across all channels to identify and merge mirror-symmetric pairs, ensuring a concise and physically sound classification. Following the selection of representative geometries, the Engineer and Student collaborate to generate trajectory evolution animations. The Engineer adapts and refines predefined code templates to construct robust visualization scripts, which are then executed by the Student to produce high-fidelity videos for the visualizations of typical trajectory evolution.
@@ 2.2 Case study
In this work, we employed keto isocytosine as a prototype system to illustrate the proposed VisU framework. As a tautomer of cytosine, one of the canonical DNA bases, the photoinduced dynamics of keto isocytosine have been extensively investigated both experimentally and theoretically. gorb1999theoretical, ha1996quantum, kwiatkowski1997density, vranken1994infrared, shukla2000investigations, bakalska2012comparative, szabla2016ultrafast, hu2017nonadiabatic, segarra2021modelling The molecular structure of keto isocytosine with atom labeling is shown in Figure 1.(d). The NAMD simulation data used in this study were taken from our previous work, zhuPrincipalComponentAnalysis2022a and relevant details are given in Supporting Information as well.

### v2

@@ Abstract
Analyzing nonadiabatic molecular dynamics trajectories traditionally heavily relies on expert intuition and visual pattern recognition, a process that is difficult to formalize. We present VisU, a vision-driven framework that leverages the complementary strengths of two state-of-the-art large language models (Doubao-Seed-1.6-Vision and DeepSeek-V3.2) to establish a “virtual research cooperation.” This operates through a “Mentor-Engineer-Student” paradigm that mimics the collaborative intelligence of a professional chemistry laboratory. Within this ecosystem, the Mentor provides physical intuition through visual reasoning, while the Engineer adaptively constructs analysis scripts, and the Student executes the pipeline and managing the data and results. VisU autonomously orchestrates a four-stage workflow comprising Preprocessing, Recursive Channel Discovery, Important-Motion Identification, and Validation/Summary. This systematic approach identifies reaction channels and key nuclear motions while generating a professional academic report at the end. By bridging visual insight with chemical expertise, VisU establishes a new paradigm for human-AI collaboration in the analysis of excited-state dynamics simulation results, significantly reducing dependence on manual interpretation and enabling more intuitive, scalable mechanistic discovery.
@@ 1 Introduction
Distilling chemically meaningful insights from the output data obtained from the trajectory-based NAMD simulations requires the integration of visual reasoning with domain-specific knowledge. Researchers routinely inspect trajectory evolutions, critical structures and dynamical animations to correlate molecular spatial patterns with photochemical concepts, such as conical intersections and reaction coordinates This tightly-coupled visual-chemical reasoning loop is essential for distinguishing nonadiabatic channels and identifying governing nuclear motions.
Recent advances in unsupervised machine learning (ML) provided valuable tools for dimensionality reduction and clustering to handle high-dimensional trajectory data. zhuUnsupervisedMachineLearning2024, zhu2022analysis, glielmo2021unsupervised, ceriottiUnsupervisedMachineLearning2019, virshup2012nonlinear, achesonAutomaticClusteringExcitedState2023, belyaev2015nonadiabatic, capano2017photophysics, delmasAutomatedSelectionNuclear2025, how2021significance, how2022dimensionality, karaDONKEYFlexibleAccurate2025, kochmanNonadiabaticMolecularDynamics2024, liAnalysisGeometricalEvolution2017, liAnalysisTrajectorySimilarity2018, linPredictionExcitedstateReaction2021, mangan2021dependence, pengAnalysisBathMotion2021a, richings2021analyzing, tavadze2018machine, zhou2020structural, liu2025automatedframeworkanalyzingstructural, tsutsumiReactionSpaceProjector2022, tsutsumiVisualizationIntrinsicReaction2018 However, these purely data-driven methods typically lack the interpretive capacity of a chemist, who perceives data patterns through a combination of visual intuition and domain-specific knowledge. Consequently, advanced automated pipelines zhuPrincipalComponentAnalysis2022a, zhuUnsupervisedMachineLearning2024, pinheiroULaMDynEnhancingExcitedstate2025, liu2025automatedframeworkanalyzingstructural still require continuous human intervention for visual validation and parameter tuning. zhuUnsupervisedMachineLearning2024 In one word, the human experts still play essential roles in such analysis protocols, due to their visiual preception and professional knowledge.
The rise of large language models (LLMs) offers new opportunities for orchestrating scientific workflows across various scientific research fields. ai4science2023impactlargelanguagemodels, jimenez2024swebench, laurent2024labbenchmeasuringcapabilitieslanguage, miret2025enabling, white2023future, jablonka202314, ramos2025review, chenSciToolAgentKnowledgeGraphDriven2025 However, conventional LLM-based approaches remain poorly aligned with the demands of NAMD analysis. li2025chemvlmexploringpowermultimodal, chenLargescaleChemicalReaction2025 Most existing working paradigms apply LLMs as either autonomous planners or passive executors. Neither role is suitable for trajectory interpretation, which requires simultaneous visual comprehension and chemical reasoning of complex spatial patterns. Critically, text-only LLMs lack access to the visual cues essential for understanding molecular geometries and reaction pathways, leaving this domain largely unexplored.
@@
By orchestrating vision-capable LLMs within this structured framework, VisU implements an end-to-end perception-to-insight pipeline that bridges the gap between raw numerical data and interpretable chemical knowledge. In this framework, visual reasoning is not merely a supplementary tool but the primary driver of the iterative process, mimicking how a scientist’s eye discerns complex patterns to refine hypotheses and validate mechanisms. This work demonstrates that embedding autonomous visual intelligence into the NAMD analysis workflow redefines how we process excited-state dynamics. At the same time, this novel research paradigm offers a scalable template for human-AI collaboration across theoretical chemistry, which enables researchers to uncover molecular mechanisms from massive datasets with unprecedented clarity and efficiency.
@@ 2.1 A Brief Description of VisU
In this section, we provide an overview of VisU, an autonomous vision-AI-powered unsupervised analysis framework.
In a traditional theoretical team, three roles perform distinct but complementary functions in a joint project: an experienced mentor supervises the overall analysis and provides expert chemical guidance, a skilled programming-engineer handles computational tool construction and processing script orchestration, and a diligent student executes computational tasks while generating intermediate results.
VisU framework adopts a “Mentor-Engineer-Student” architecture, in which these three roles are designed faithfully mirroring the division of labor in a real nonadiabatic dynamics research group.
  • The Mentor: Powered by the Doubao-Seed-1.6-Vision LLM, the Mentor integrates deep chemical insights with advanced vision-reasoning capabilities to oversee the analytical direction.
  • The Engineer: Built upon the DeepSeek-V3.2 LLM equipped with specialized computational tools and coding templates, the Engineer demonstrates robust programming and debugging proficiency.
  • The Student: Responsible for data management, script execution, and visualization, the Student serves as the operational hub, providing feedback to the Mentor and Engineer.
Following this human-mimetic paradigm, VisU operationalizes NAMD analysis into a fully automated, four-stage end-to-end protocol:
@@
To demonstrate its potential, the current implementation of VisU builds upon the validated framework of previous work. zhuPrincipalComponentAnalysis2022a That work established a hierarchical analysis protocol for characterizing ring deformations in atomic systems via a combination of principal component analysis (PCA) wold1987principal, abdi2010principal and various clustering algorithmsester1996density, schubert2017dbscan, kriegel2011density, ward1963hierarchical, kaufman2009finding. Future versions will expand the framework’s applicability to a broader spectrum of molecular architectures and complex photophysical processes.
@@
[H] [Uncaptioned image] Workflow of VisU, an autonomous vision-AI framework for hierarchical analysis of nonadiabatic dynamics. The VisU system comprises the three-role paradigm and four functional modules: (a) “Mentor-Engineer-Student” architecture; (b) Preprocessing module autonomously parses raw NAMD outputs and generates standardized descriptor sets; (c) Recursive Channel Discovery module iteratively applies PCA-guided clustering to uncover chemically distinct decay channels; (d) Important-Motion Identification module analyzes PCA projections and visualizes dominant nuclear displacements for each channel; (e) Validation and Mechanistic Summary module selects representative structures, performs final quality checks, and generates concise mechanistic interpretations.
@@ 2.1.1 Data Preprocessing
VisU accepts two input formats: (i) three mandatory XYZ files, that is, ref.xyz (reference geometry of ground-state minimum), hopping.xyz (all hopping geometries from NAMD trajectories), and init.xyz (corresponding initial geometries), or (ii) a single ZIP archive containing raw simulation outputs from any NAMD package. Although both approaches are allowed, the direct submission of the three XYZ files is highly recommended for routine analyses, as it saves tokens and eliminates any remaining risk of extraction errors.
@@
In general, different NAMD packages produce distinct output formats and directory structures. Conventional LLM-based workflows lack direct access to file contents and system details, preventing them from reliably interpreting such heterogeneous file systems during preprocessing. Without knowledge of the file contents or structures, these pipelines typically rely on multistep trial-and-error workflows, iteratively attempting to interpret outputs, generate scripts, execute them, and refine the logic based on error messages. This process not only incurs high computational costs and carries a substantial risk of non-convergence but also requires extensive elaborate instruction frameworks to handle diverse scenarios.
@@
To ensure broad compatibility across the heterogeneous and often inconsistent output file formats generated by different NAMD simulation packages, we deliberately avoid allowing LLMs to generate a complete preprocessing script from scratch. Instead, the framework provides a curated library of expert-validated Python templates for structure extraction, format sanitization, and data integrity checks. With the support of these templates, the Engineer constructs the full preprocessing script by assembling and adapting predefined components rather than writing arbitrary code. This design substantially improves both flexibility and reliability in data preprocessing, surpassing conventional LangChain-style agents that are often constrained by static tool graphs or rigid, hard-coded workflows. Following script generation, the Student executes the preprocessing codes, generates all required files, and returns execution feedback. If execution fails, the complete traceback information is automatically fed back to the LLM, enabling iterative self-repair and refinement of the preprocessing operations until valid output files are produced. Overall, this template-grounded and self-repairing strategy establishes a more effective paradigm for automated scientific code generation and robust preprocessing of heterogeneous simulation outputs. Crucially, the tools and templates developed for the Engineer are designed to be general-purpose and system-agnostic, rather than being tailored to specific case studies. More details are provided in the SI.
@@
Following preprocessing, VisU proceeds to systematic structure analysis and descriptor generation through the close cooperation of the Mentor (built on Doubou-Seed-1.6-Vision) and Student. The core of this stage is the introduction of a dedicated vision-language chemistry Mentor, a large multimodal model extensively prompted with expert-level chemical knowledge. The workflow starts with the Student generating a high-quality 2D molecular diagram from the optimized reference geometry in ref.xyz.
@@
As the currently supported analysis workflow in VisU primarily targets the geometric evolution of atomic ring systems in NAMD simulations, the following discussion focuses on such cases. This scope is sufficient to illustrate the effectiveness of the vision-powered “Mentor-Engineer-Student” cooperative strategy for NAMD analysis, while extensions to more general molecular topologies are left for future work.
Previous works zhuPrincipalComponentAnalysis2022a, liu2025automatedframeworkanalyzingstructural demonstrated that redundant internal coordinates (RICs) pulay1992geometry, peng1996using provide an appropriate molecular descriptor for NAMD analysis and are therefore adopted here. RICs consist of bond lengths (R), bond angles (A), and dihedral angles (D), which span different numerical ranges and encode distinct physical aspects of molecular structure. For systems containing atomic rings, RICs associated with the ring framework and those involving terminal groups play fundamentally different roles during nonadiabatic dynamics. Accordingly, within the workflow, the Mentor instructs the Student to classify atoms into ring (ring) and end-group (eg) subsets. The Student then constructs six descriptor groups, $R_{\mathrm{ring}}$, $A_{\mathrm{ring}}$, $D_{\mathrm{ring}}$, $R_{\mathrm{eg}}$, $A_{\mathrm{eg}}$, and $D_{\mathrm{eg}}$, which are subsequently used for unsupervised analysis. To avoid overemphasizing large-amplitude hydrogen motions, hydrogen atoms are excluded, and the analysis focuses exclusively on the molecular scaffold.
@@ 2.1.2 Recursive Channel Discovery
The selection of an appropriate descriptor subset follows a standardized visual screening procedure. The Student generates six low-dimensional projections corresponding to all descriptor groups and submits them to the Mentor for inspection. Based on the separability, compactness, and overall structure of the data distributions, the Mentor identifies the reduced spaces that exhibit the most distinguishable channel patterns and selects the corresponding reduced space for subsequent clustering analysis.
@@ 2.1.3 Important-Motion Identification
The identification of these critical motions is performed through a tight coordination between the Mentor and the Student. For each identified channel, the Student assembles the initial and hopping geometries, executes PCA, and generates a comprehensive package. This package includes 2D projection plots, heatmaps of the principal components, and three-view (front, side, and top) renderings of both the reference and averaged hopping structures. Under the guidance of its vision-reasoning capabilities, the Mentor evaluates these multi-faceted figures to pinpoint the active coordinates and leading molecular motions characteristic of the channel. This Mentor-Student joint procedure effectively bridges the gap between unsupervised data clustering and human-interpretable chemical mechanisms, establishing a fully automated NAMD analysis protocol that eliminates the requirement for manual validation.
@@ 2.1.4 Validation and Mechanistic Summary
For each reaction channel, the Student randomly samples a representative set of hopping geometries (typically ten) and prepares their three-view (front, side, and top) renderings for the Mentor. Upon visual evaluation, the Mentor identifies the most characteristic geometry based on previously established analysis results. Simultaneously, the Mentor inspects representative geometries across all channels to identify and merge mirror-symmetric pairs, ensuring a concise and physically sound classification. Following the selection of representative geometries, the Engineer and Student collaborate to generate trajectory evolution animations. The Engineer adapts and refines predefined code templates to construct robust visualization scripts, which are then executed by the Student to produce high-fidelity videos for the visualizations of typical trajectory evolution.
@@ 2.2 Case study
In this work, we employed keto isocytosine as a prototype system to illustrate the proposed VisU framework. As a tautomer of cytosine, one of the canonical DNA bases, the photoinduced dynamics of keto isocytosine was extensively investigated both experimentally and theoretically. gorb1999theoretical, ha1996quantum, kwiatkowski1997density, vranken1994infrared, shukla2000investigations, bakalska2012comparative, szabla2016ultrafast, hu2017nonadiabatic, segarra2021modelling The molecular structure of keto isocytosine with atom labeling is shown in Figure 1.(d). The NAMD simulation data used in this study were taken from our previous work, zhuPrincipalComponentAnalysis2022a and relevant details are given in SI as well.